OpenStudy (anonymous):

Find the product. 1/10 • 4/10

OpenStudy (anonymous):

i want you to try it

OpenStudy (anonymous):

\[\frac{ a }{ b }*\frac{ c }{ d }=\frac{ ac }{ b }\]

OpenStudy (anonymous):

\[\frac{ ac }{ bd }\]

OpenStudy (anonymous):

\[\frac{ 1*4 }{ 10*10 }\] \[\frac{ 4 }{ 100 }\] \[\frac{ 1 }{ 25 }\]